# Noise2Ghost: Self-supervised deep convolutional reconstruction for ghost imaging

### v1

@@ III-B Synthetic data reconstructions across various noise levels
Figure 3: Synthetic data GI reconstructions of the chromosomes phantom, with 5$\times$ compression ratio, and varying noise levels. The reconstruction algorithm performance is compared for maximum emitted photons per pixel per realization intensities in the range of [$10^{0}$, $10^{4}$]: (3a) reconstruction peak signal-to-noise ratio (PSNR) and structural similarity index (SSIM) against the ground truth, where higher values indicate higher reconstruction quality; and (3b) computed bandwidths in terms of the Nyquist frequency and corresponding resolutions in pixels of the images.
We present three different study cases: In the first case (Fig. 3), we present the performance of the above-mentioned different algorithms (LS, TV, GIDC, INR, and N2G) across multiple noise levels (i.e. photons per pixel) using only 2000 realizations, corresponding to 20% of the total number of pixels. The simulated maximum emitted photons per pixel per realization (i.e. constant $C$) intensities span the range of $[10^{0},10^{4}]$. For each noise level, we generated five different sets of realizations. The same evaluation metrics (e.g. mean error against the phantom, etc) are independently computed for each of the five reconstructions. The presented results indicate the average and standard deviation of the computed metrics.
For the TV, GIDC, INR, and N2G reconstructions, the regularization term $\lambda$ was chosen by cross-validation, by putting aside a set of realizations ($10\%$ of the total) and using them to assess the best value over a wide range [27]. For TV, we use a $\lambda\in[10^{-5},10^{-1}]$, while for GIDC, INR, and N2G we used a $\lambda\in[10^{-7},10^{-4}]$. To improve the procedure’s robustness, we ran it three times with different cross-validation sets for each run and averaged the cross-validation loss values between those different runs. For the INR reconstructions, we used a 3-layer feed-forward NN with 512 neurons per layer, sinusoidal activation functions [15], and 64 Fourier embeddings [28]. Regarding specifically GIDC and N2G, we used a U-net [11] with 24 features and 3 decomposition levels ($\sim$ 276k parameters). We fitted the U-net weights with 3000 epochs of the Adam optimizer [29] with a learning rate of $2\times 10^{-4}$ and weight decay of $1\times 10^{-2}$. We selected the model weights of the epoch where the cross-validation loss function is minimized. For N2G, we used 4 partitions and 6 permutations ($K=4$, and $P=6$).
@@
Figure 4: Dose requirements for the chromosomes phantom, with 5$\times$ compression ratio, and varying noise levels. The reconstruction algorithms performance is compared against PB, for maximum emitted photons per pixel per realization intensities in the range of [$10^{0}$, $10^{4}$]: (4a) total required dose for each algorithm to obtain equivalent PSNR and SSIM as a PB acquisition; (4b) same plot as (4a) for the average maximum pixel dose per illumination; and (4c) for N2G.
On the left side of Fig. 3a, we present the peak signal-to-noise ratio (PSNR) of the reconstructions against the phantom. It is evident that N2G consistently achieves superior reconstructions, with PSNRs surpassing those of competing methods by 1–2 dB, highlighting its robustness across diverse imaging conditions. Nonetheless, its advantage seems to diminish under low-photon scenarios. While counterintuitive, this is easily explained by the fact that the lower frequencies are less affected by Poisson noise. PSNR takes all frequencies into account, resulting in a bias towards lower frequencies with low-photon counts. The SSIM, by construction, is instead less affected by lower frequencies. For this reason, on the right side of Fig. 3a, we see that N2G has a growing lead in SSIM for lower photon counts. We also observe that for high photon numbers (equating to low noise), the compression ratio becomes the limiting factor for reconstruction accuracy, and both PSNR and SSIM reach a plateau. From Fig. 3b, we have instead a perspective on how the resolution is degraded with increasing noise levels. We observe that N2G better preserves the reconstruction resolution, by showing a drop-off at lower photon counts than other methods.
Finally, we explore dose considerations by comparing GI with PB and assess N2G’s performance against LS, TV, GIDC, and INR. In particular, we evaluate the total sample and average maximum pixel dose necessary to achieve a GI image with PSNR or SSIM quality comparable to a PB scan. From figure Fig. 4a, we observe that in terms of the total dose, PB always outperforms GI at the chosen compression ratio, regardless of the method used, although N2G narrows the difference. Concerning the average maximum pixel dose per exposure (Fig. 4b), GI generally exceeds PB’s performance when the main noise source is detection noise (except when LS is used). However, N2G still decreases dose requirements compared to the other GI reconstruction techniques and widens the region where GI is preferable over PB. Figure 4c clearly shows this trend by plotting the required dose ratio from LS, TV, GIDC, and INR to obtain the same performance as N2G.
III-C In-depth analysis of noisy reconstructions
@@
Figure 5: GI reconstructions of the chromosomes phantom, with 10$\times$ compression ratio and moderate Poisson noise: Mean noise fluctuations $\sim 24.5\%$ of the mean clean bucket fluctuations. From (5a) to (5e) we show the reconstructions of (5f) with LS, TV, GIDC, INR and N2G respectively. In (5g) and in (5h) we present the Fourier ring correlation against (5f) and various performance metrics of each reconstruction, respectively.
@@
We present the results for (a) in Fig. 5. We produced 1000 realizations, resulting in a 10$\times$ compression ratio. The applied Poisson noise of $10^{2}$ maximum emitted photons per pixel per realization (i.e. constant $C=10^{2}$) resulted in fluctuations equal to $24.5\%$ of the mean clean bucket fluctuations. Comparing reconstructions from Figs. 5a to 5e against Fig. 5f, we note that except for the LS reconstruction, all other algorithms, including N2G, can correctly reconstruct the shape of the phantom. The only visible difference is that N2G’s reconstruction presents much less high-frequency noise. This results in the N2G reconstruction having the highest resolution (in pixels), PSNR, and SSIM values, and the lowest MSE among all reconstructions. These results are summarized in Fig. 5h.
@@
Figure 6: GI reconstructions of the chromosomes phantom, with 3$\times$ compression ratio and very high Poisson noise: Mean noise fluctuations $\sim 164.44\%$ of the mean clean bucket fluctuations. From (6a) to (6e) we show the reconstructions of (6f) with LS, TV, GIDC, INR and N2G respectively. In (6g) and in (6h) we present the Fourier ring correlation against (6f) and various performance metrics of each reconstruction, respectively.
We present the results of (b) in Fig. 6. Here, we produced $3$ realizations, resulting in a 3$\times$ compression ratio. The applied Poisson noise of 2 maximum emitted photons per pixel per realization (i.e. constant $C=2$) resulted in fluctuations equal to $164.44\%$ of the mean clean bucket fluctuations. All the reconstructed images are noticeably affected by noise with a larger bandwidth than in the previous examples. N2G can suppress high-frequency noise, but cannot completely correct lower-frequency noise. Despite that, by looking at the FRC in Fig. 6g, we notice that it preserves much more bandwidth than the other algorithms and presents a much better resolution. This again results in the N2G reconstruction having the highest PSNR and SSIM values, and the lowest MSE among all reconstructions. These results are summarized in Fig. 6h.
III-D X-ray fluorescence data reconstructions
Finally, we compare the reconstruction performance of N2G on real GI data acquired at the beamline ID19 of the ESRF — The European Synchrotron [30]. The considered dataset consists of $42\times 87$ pixels images, with $24\mu$m pixel size, and 896 realizations, where the buckets are the emitted XRF signals from a sample composed of a glass capillary and three wires. Two of the three wires are made of Cu, and they will be the focus of this reconstruction. We used the signal from their $K_{\alpha}$ emission line. At each realization, we exposed the sample to a 26 keV incoming X-ray beam for 0.1 seconds and acquired each realization 32 times. The accumulated 3.2-second exposures serve as high-SNR buckets. Reducing the number of accumulated exposures artificially reduces the collected number of photons (photon flux), thus simulating lower dose depositions. The XRF signals are collected with a single-pixel hyper-spectral detector, whose spectral output is discretized into 4096 bins of 150 eV energy steps. Each XRF peak has a Lorentzian shape around the mean energy $E_{e,l}$. This means that we observe photon counts in adjacent bins around the expected mean signal energy, which decrease $\sim\gamma_{e,l}/((E_{e,l}-E)^{2}+\gamma_{e,l}^{2})$ for bins at energy $E$ further away from it, where $\gamma_{e,l}$ is a constant that depends on the emission line. These neighboring bins are either summed or the area under their curve is fitted to increase the signal’s SNR. Therefore, we can simulate further dose reductions by summing fewer and lower-count bins.
@@ III-D X-ray fluorescence data reconstructions
d PSNR and SSIM of the GI reconstructions.
Figure 7: Real data GI reconstructions of XRF-GI data from [30], using TV, GIDC, and N2G: (7a) presents the reconstructions at different photon counts (and corresponding noise levels); (7b) and (7c) present an x ray transmission image of the sample at 26 keV and the segmentation of the Cu wires, respectively; and (7d) a table of the peak signal-to-noise ratio (PSNR) and structural similarity index (SSIM) of each reconstruction against their high-SNR versions in the top row of (7a).
We present the results of the XRF-GI reconstructions in Fig. 7. Figure 7b presents an X-ray transmission image (radiograph) of the sample, while Fig. 7c indicates the regions that give rise to the XRF signal of interest. As mentioned above, we consider the 0.1-second exposures as the “100% dose” exposures, while the 3.2-second accumulated images serve as ground truth (high-quality reference images). The other two images in Fig. 7a were produced using the five highest-counts and the single highest-counts XRF bins, which provide an average of 12.9% and 2.6% of the reference dose per realization, respectively. These two reduced-flux datasets could thus correspond to 12.9 ms and 2.6 ms exposures per realization, respectively. The table from Fig. 7d presents the estimated image quality degradation with increasing noise levels, according to PSNR and SSIM. Each result is compared against their own method-related “reference” reconstruction because we lack a PB scan of the object, which would serve as shared ground truth. Despite that, N2G provides the highest noise reduction according to PSNR and SSIM across all the considered reconstruction methods.
IV Conclusions & Outlook
@@
The above-mentioned aspects also mean that our method could enable operating GI scans at $\sim$ 1.5 – 2 shorter dwell time, resulting in faster GI acquisitions. This aspect is crucial for real case scenarios in two alternative ways: Either increasing acquisition stability or the ability to acquire more realizations for the same dose. In the first case, when paired with the reduced number of GI realizations compared to PB acquisitions, we further reduce the total acquisition times of GI acquisitions, strongly improving the stability of our GI scans. Moreover, compared to PB scans, where the sample moves across the field of view, in a GI acquisition, the sample is fixed [30]. This further consolidates GI as a very attractive technique for reducing sample drifts and positioning errors in nano-scale imaging. Alternatively, the reduced exposure time for each GI realization would also enable more realizations to be acquired per a fixed acquisition time budget. This, in turn, is the only way to reduce the missing-realization noise, which is the ultimate limit in GI reconstruction quality, as seen earlier in the text.
N2G focuses on noise reduction from the acquired measurements. It delegates dealing with the missing realization artifacts to the generalization power of the underlying model (NN) and the rather simplistic prior knowledge given by the TV term. Future work could tackle this problem with a combination of a few possible approaches, including using stronger or more complex priors than TV (e.g., multi-level undecimated wavelet minimization) and multi-channel/multi-modal information from correlated signals (e.g., the transmission signal in [16]), as demonstrated in computed tomography [31].
-A Metrics
We give the mathematical definition of the metrics used in section III. For two given signals $x_{1}$ and $x_{2}$, MSE, PSNR, and SSIM can be computed as follows:
$\mbox{MSE}(x_{1},x_{2})=\lVert x_{1}-x_{2}\rVert_{2}^{2}$ (14a)
$\mbox{PSNR}(x_{1},x_{2})=20\,log_{10}\lVert x_{1}\rVert_{\infty}-10\,log_{10}\,\mbox{MSE}(x_{1},x_{2})\;$ (14b)
$\mbox{SSIM}(x_{1},x_{2})=\frac{(2\mu_{1}\mu_{2}+c_{1})(2\sigma_{12}+c_{2})}{(\mu_{1}^{2}+\mu_{2}^{2}+c_{1})(\sigma_{1}^{2}+\sigma_{2}^{2}+c_{2})}\;$ (14c)
where $\lVert x_{1}\rVert_{\infty}=max(x_{1})$, $\mu$ and $\sigma$ are the mean value and the standard deviation respectively of a given signal, the $c$ constants are used to stabilize the fraction, and the SSIM is computed on $7\times 7$ pixel windows. From [26], the FRC is defined as:
$\mbox{FRC}_{x_{1},x_{2}}(r_{i})=\frac{\sum_{r\in r_{i}}\mathcal{F}_{x_{1}}(r)\cdot\mathcal{F}_{x_{2}}^{*}(r)}{\sqrt{\sum_{r\in r_{i}}\mathcal{F}_{x_{1}}^{2}(r)\cdot\sum_{r\in r_{i}}\mathcal{F}_{x_{2}}^{2}(r)}}$ (15)
where $\mathcal{F}_{x}(\cdot)$ is the Fourier transform of a signal $x$, and $r\in r_{i}$ is the set of all the pixels at distance $r_{i}$ from the zero frequency in Fourier space, resulting in $\sum_{r\in r_{i}}$ being an azimuthal integration at the radius $r_{i}$.
-B Derivation of the noise model
Since the proposed method is agnostic of the underlying imaging technique, we first develop an independent evaluation framework from the technique and noise source. We use the guiding example of XRF imaging to derive the just-mentioned generic noise model. For PB acquisitions and a given element $e$ in a compound $cmp$ and the corresponding XRF emission line $l$, the clean observed signal in the XRF detector is (slightly adapted from Eq. (10) of [32]):
$\Phi_{e,l}(E_{0})=\phi_{0}(E_{0})t\chi_{e}\sigma_{e,l}(E_{0})\rho_{cmp}\nu G_{d}A_{e,l}(E_{0})$ (16a)
$\nu=p^{2}\tau$ (16b)
$G_{d}=\frac{S_{d}}{4\pi D_{d}^{2}}$ (16c)
@@
where $\phi_{0}(E_{0})$ is the probing beam’s photon flux at the energy $E_{0}$, $t$ the dwell time (exposure time per pixel), $\rho_{c}$ is the local compound density, $\chi_{e}$ is the local relative concentration of the said element $e$ in the compound $cmp$, $\sigma_{e,l}(E_{0})$ is the XRF production cross-section, $\nu$ is the observed volume, where $p$ is the beam waist (pixel size) and $\tau$ the object thickness, $G_{d}$ is a geometric factor that accounts for the detector surface $S_{d}$ and distance $D_{d}$, and finally $A_{e,l}(E_{0})$ is the combined attenuation of the incoming and emitted photons across the sample. The total delivered dose per pixel is $\Phi_{0}(E_{0})=\phi_{0}(E_{0})t$. Equation 16 can be easily modified to account for other emission signals like Compton radiation. The corresponding detected signal is $\Phi_{\alpha}(E_{0})$, where $\alpha$ is the observation angle, is obtained by changing the production cross section $\sigma_{e,l}(E_{0})$ to the Compton equivalent, and by removing the elemental concentrations $\chi_{e}$ (as $\sum_{e}\chi_{e}=1$).
The observed signal $y$ on the detector is:
$y=\mathcal{P}\left(\Phi_{e,l}(E_{0})\right)+\mathcal{N}_{\mu_{d}=0,\sigma_{d}^{2}}$ (17)
where $\mathcal{P}$ is the photon counting (Poisson – shot) noise of the detection process, and $\mathcal{N}$ is the additive white (Gaussian) noise from the instrument electronics, with mean value $\mu_{d}=0$ and a given variance $\sigma_{d}^{2}$. Per Eq. 1, $b=\Phi_{e,l}(E_{0})$ and $\epsilon=(\mathcal{P}\left(\Phi_{e,l}(E_{0})\right)-\Phi_{e,l}(E_{0}))+\mathcal{N}_{\sigma_{d}^{2}}$. In x-ray emission detection, it is safe to assume $\sigma_{d}^{2}=0$. Thus, we are mostly confronted with Poisson noise, which is mainly modulated by the quantities $\Phi_{0}(E_{0})$, $\rho_{c}$, and $\chi_{e}$.
In XRF imaging, we aim to reconstruct the local relative concentrations $\chi_{e}$ (or actual concentrations $\chi_{e}\rho_{c}$), while in Compton imaging, we aim at the compound densities $\rho_{c}$. Thus, they correspond to the vector elements $x_{i}$ from Eq. 1. This means that the other factors in Eq. 16a can either be considered constant or measured during the acquisitions, and they will contribute as constants in our noise description. Therefore, e.g. for XRF imaging, we can simplify Eqs. 16a and 17 to account for each measured point $j$ as follows:
$b_{m}=\Phi_{e,l,m}(E_{0})=\chi_{e,m}c_{e,l}(E_{0})\Phi_{0}(E_{0})$ (18a)
$y_{m}=\mathcal{P}\left(\chi_{e,m}c_{e,l}(E_{0})\Phi_{0}(E_{0})\right).$ (18b)
Thus, the noise is modulated by the value of $C=c_{e,l}(E_{0})\Phi_{0}(E_{0})$, giving $b=\chi_{e}C$ and $y=\mathcal{P}(\chi_{e}C)$. The constant $C$ represents the emitted photons per pixel $j$ for a 100% local element $e$ concentration.
For GI and a homogeneous incoming beam, we modify Eq. 18 as follows:
$b_{m}=\Phi_{e,l,m}(E_{0})=\textstyle\sum_{n}^{N}w_{mn}\chi_{e,n}c_{e,l}(E_{0})\Phi_{0}(E_{0})$ (19a)
$y_{m}=\mathcal{P}\left(\textstyle\sum_{n}^{N}w_{mn}\chi_{e,n}c_{e,l}(E_{0})\Phi_{0}(E_{0})\right)$ (19b)
where $w_{mn}$ are the mask transmission intensities from the projection matrix $W$ in Eq. 1. If $w_{mn}\in[0,1]$, the constant $C$ represents the maximum emitted photons per pixel. As expected, from Eq. 19 we observe that $W_{PB}=I$, for PB acquisitions. By absorbing the $c_{e,l}(E_{0})\Phi_{0}(E_{0})$ into the the same constant $C$ as for PB, and normalizing data by $C$, we obtain Eq. 13.
-C Deposited dose & acquisition time considerations
Equations 18 and 19 offer a way to compare the performance of PB and GI approaches on simulated data, especially when comparing the respective simulated acquisition time or delivered dose to the sample. For PB acquisitions, we define:
@@
while for GI acquisitions:
[…]
for a given pixel $n$, where $w_{\cdot n}$ indicates the $n$-th column of the matrix $W$. If the matrix elements $w_{mn}$ span $[0,1]$, then $sup(W)=1$ and thus we can assume $\lVert w_{\cdot n}\rVert_{\infty}=1$.
Given the ground truth $\hat{x}$, a GI reconstruction $x_{GI}$ obtained from synthetic data generated with a value $C_{GI}$, and a metric $\mathcal{M}(\cdot)$ of choice in the set $\{\mbox{PSNR}(\hat{x};\cdot),\mbox{SSIM}(\hat{x};\cdot),\ldots\}$, we obtain the $\hat{C}_{PB}$ value of a PB acquisition $x_{PB}=y_{PB}=\mathcal{P}\left(\hat{x}C_{PB}\right)$ that exhibits the closest performance to $x_{GI}$ as:
$\hat{C}_{PB}=\operatorname*{\arg\!\min}_{C_{PB}}\left\{\mathcal{M}(x_{GI})-\mathcal{M}(\mathcal{P}\left(\hat{x}C_{PB}\right))\right\}.$ (22)
Finally, we can compute acquisition time or deposited dose ratios between different techniques, for equivalent image quality, by computing the ratio between the corresponding quantities in Eqs. 20 and 21, with the corresponding $C$ values from Eq. 22.

### v2

@@ 3.2 Synthetic data reconstructions across various noise levels
[…]
We present three different study cases: In the first case (Fig.˜3), we present the performance of the above-mentioned different algorithms (LS, TV, GIDC, INR, and N2G) across multiple noise levels (i.e., photons per pixel) using only 2000 realizations, corresponding to 20% of the total number of pixels. The simulated maximum emitted photons per pixel per realization (i.e., constant $C$) intensities span the range of $10^{0},10^{4}]$. The lower margin means that each pixel is illuminated on average by around one photon. The upper margin was empirically chosen well above the threshold where noise is not more the limiting factor of the reconstruction quality (i.e., it becomes the missing realizations). To add some statistical significance to our tests, we generated five different sets of realizations for each noise level. We then independently reconstruct these five sets and compute their individual evaluation metrics (e.g., mean error against the phantom, etc). The presented results report the average and standard deviation of the said computed metrics for each noise level.
For the TV, GIDC, INR, and N2G reconstructions, the regularization weight $\lambda$ was chosen by cross-validation, by putting aside a set of realizations ($10\%$ of the total) and using them to assess the best value over a wide range [21]. For TV, we use a $\lambda\in[10^{-5},10^{-1}]$, while for GIDC, INR, and N2G we used a $\lambda\in[10^{-7},10^{-4}]$. This difference in range is due to the different data normalization used by the different algorithm implementations. As regularization $R$ for GIDC, INR and N2G, we also used the TV term. Although stronger and more advanced priors could have been used, we wanted a direct comparison against the well-known and used TV-min-based variational regularization. Moreover, the TV term has very simple and widely understood implementation and behavior. To improve the procedure’s robustness, we ran it three times with different cross-validation sets for each run and averaged the cross-validation loss values between those different runs. For the INR reconstructions, we used a 3-layer feed-forward NN with 512 neurons per layer, sinusoidal activation functions [18], and 256 Fourier embeddings [26]. Regarding specifically GIDC and N2G, we used a U-net [14] with 24 features and 3 decomposition levels ($\sim$ 276k parameters). We learned the U-net and INR weights with 3000 and 7000 epochs of the Adam optimizer [27] respectively, with a learning rate of $2\times 10^{-4}$ and weight decay of $1\times 10^{-2}$ for both architectures. We selected the model weights corresponding to the epoch in which the cross-validation loss function is minimized. For N2G, we used 4 partitions and 6 permutations ($K=4$, and $P=6$). These hyperparameter values were also selected cross-validation over a grid search.
@@
[…]
On the left side of Fig.˜3a, we present the peak signal-to-noise ratio (PSNR) of the reconstructions against the phantom. It is evident that N2G consistently achieves superior reconstructions, with PSNRs surpassing those of competing methods by 1–2 dB, highlighting its robustness across diverse imaging conditions. Nonetheless, its advantage seems to diminish under low-photon scenarios. While counterintuitive, this is easily explained by the fact that the lower frequencies are less affected by Poisson noise. PSNR takes all frequencies into account, resulting in a bias towards lower frequencies with low photon counts. The SSIM, by construction, is instead less affected by lower frequencies. For this reason, on the right side of Fig.˜3a, we see that N2G has a growing lead in SSIM for lower photon counts. We also observe that for high photon numbers (equating to low noise), the compression ratio becomes the limiting factor for reconstruction accuracy, and both PSNR and SSIM reach a plateau. From Fig.˜3b, we have instead a perspective on how the resolution is degraded with increasing noise levels. We observe that N2G better preserves the reconstruction resolution, by showing a drop-off at lower photon counts than other methods.
Finally, we explore dose considerations by comparing GI with PB and assess N2G’s performance against LS, TV, GIDC, and INR. In particular, we evaluate the total sample and average maximum pixel dose necessary to achieve a GI image with PSNR or SSIM quality comparable to a PB scan. From figure Fig.˜4a, we observe that in terms of total dose, PB always outperforms GI in the chosen compression ratio, regardless of the method used, although N2G narrows the difference. Concerning the average maximum pixel dose per exposure (Fig.˜4b), GI generally exceeds PB’s performance when the main noise source is detection noise (except when LS is used). However, N2G still decreases dose requirements compared to the other GI reconstruction techniques and widens the region where GI is preferable over PB. Figure˜4c clearly shows this trend by plotting the required dose ratio from LS, TV, GIDC, and INR to obtain the same performance as N2G.
3.3 In-depth analysis of noisy reconstructions
@@
[…]
@@
[…]
@@
[…]
As a final remark, we observe that in both the presented cases, N2G tends to provide even smoother backgrounds than the phantom itself. The overly smoothed features are usually high-frequency and low-signal features, which are much more overwhelmed by noise than the high-signal parts of the image. As we discussed in Sec.˜2, the suppression of high-frequency noise can result in oversmoothing in these cases.
3.4 X-ray fluorescence data reconstructions
[…]
@@ 3.4 X-ray fluorescence data reconstructions
Dose
Avg. photons 32000% 100% 12.9% 2.6%
Bucket 2.07e+05 6.47e+03 8.38e+02 1.69e+02
Bucket & pixel 3.48 0.109 0.0141 0.00283
d Photon counts of each acquisition.
@@
e PSNR and SSIM of the GI reconstructions.
Figure 7: Real data XRF-GI reconstructions: (7a) reconstructions at different photon counts (and corresponding noise levels); (7b) and (7c) an x ray transmission image of the sample at 26 keV and the segmentation of the Cu wires, respectively; (7d) the photoncounts of each acquisition; and (7e) PSNR and SSIM of each reconstruction against their high-SNR versions in the top row of (7a).
We present the results of the XRF-GI reconstructions in Fig.˜7. Figure˜7b presents an X-ray transmission image (radiograph) of the sample, while Fig.˜7c indicates the regions that give rise to the XRF signal of interest. As mentioned above, we consider the 0.1-second exposures as the “100% dose” exposures, while the 3.2-second accumulated images serve as ground truth (high-quality reference images). The other two images in Fig.˜7a were produced using the five highest-counts and the single highest-counts XRF bins, which provide an average of 12.9% and 2.6% of the reference dose per realization, respectively. These two reduced-flux datasets could thus correspond to 12.9 ms and 2.6 ms exposures per realization, respectively. The corresponding average photons per bucket and per bucket & pixel can be found in Fig.˜7d. The table from Fig.˜7e presents the estimated image quality degradation with increasing noise levels, according to PSNR and SSIM. Each result is compared against its own method-related “reference” reconstruction because we lack a PB scan of the object, which would serve as shared ground truth. Despite that, N2G provides the highest noise reduction according to PSNR and SSIM across all the reconstruction methods considered.
4 Conclusions & Outlook
@@
The above-mentioned aspects also mean that our method could enable operating GI scans at $\sim$ 1.5 – 2 shorter dwell time, resulting in faster GI acquisitions. This aspect is crucial for real case scenarios in two alternative ways: Either increasing acquisition stability or the ability to acquire more realizations for the same dose. In the first case, when paired with the reduced number of GI realizations compared to PB acquisitions, we further reduce the total acquisition times of GI acquisitions, strongly mitigating the impact of sample drifts in our GI scans. Moreover, compared to PB scans, where the sample moves across the field of view, in a GI acquisition, the sample is fixed [21]. This further consolidates GI as a very attractive technique for reducing sample drifts and positioning errors in nano-scale imaging. Alternatively, the reduced exposure time for each GI realization would also enable more realizations to be acquired within a fixed acquisition time budget. This, in turn, is the only way to reduce the missing-realization noise, which is the ultimate limit in GI reconstruction quality, as seen earlier in the text.
N2G focuses on the reduction of noise from acquired measurements. It delegates dealing with the missing realization artifacts to the generalization power of the underlying model (NN) and the rather simplistic prior knowledge given by the TV term. Future work could tackle this problem with a combination of a few possible approaches, including using stronger or more complex priors than TV (e.g., multi-level undecimated wavelet minimization) and multi-channel/multi-modal information from correlated signals (e.g., the transmission signal in [10]), as demonstrated in computed tomography [29].
@@
Funding Placeholder text for the funding, as required by the submission system.
@@
Data Availability Statement Data underlying the experimental results are available at [28].
@@
Supplemental document See Supplement 1 for supporting content.